If the sides of the parallelogram and an included angle are 6, 10 and 100 degrees respectively, find the length of the larger diagonal.


1
Expert's answer
2021-05-13T03:11:53-0400

let a=6, b=10, diagonal =d

"d^2=a^2+b^2-2abcos100^o=6^2+10^2-2\times6\times 10 cos100^o"

"d^2=36+100+20.76=156.76"

"d=\sqrt{156.76}=12.52"